OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

f(x)=8/x and g(x)=8/x

OpenStudy (anonymous):

@Cydney_Morgan @ObeySoph▲ @Whitemonsterbunny17

OpenStudy (anonymous):

@bibby